Cougar Company is owned equally by Cat Stevens and a partnership that is owned equally by his father and two unrelated individuals. Cat and the partnership each own 3,000 shares in the company. Cat wants to reduce his ownership in the company, and it is decided that the company will redeem 1,500 of his shares for $25,000 per share. Cat’s income tax basis in each share is $5,000. What are the income tax consequences to Cat as a result of the stock redemption, assuming the company has earnings and profits of $10 million? (Enter your answers in dollars and not in millions of dollars.)
Answer #1

Share of his father in shares of company = 3,000 X 1/3 = 1,000

These are deemed to be owned by Cat

Deemed ownership of cat before redemption = 4,000/6,000 X 100 = 66.67%

Deemed ownership after redemption = (4,000 - 1,500) / (6,000 - 1,500) X 100 = 55.56%

Percentage of ownership relative to before redemption = 55.56%/66.67% X 100 = 83.33%

Ownership did not decline by more than 80% of what it was before redemption.

Redemption is treated as dividend payment.

Total payment = 25,000 X 1,500 = 37,500,000

Distribution to the extent of 10 million is treated as dividend, 15 million is treated as return of capital (stock basis), and balance 10.5 million is capital gain.